OpenStudy (anonymous):

Find the radius of convergence of the series \[\sum_{n=0}^{\infty} x ^{2n+1} / (2n+1)!!\] and show that its sum f solves the equation f'(x)=1+xf(x) for all x in the interval of convergence

OpenStudy (anonymous):

is this \[\sum\frac{x^{n+1}}{(n+1)!}\]?

OpenStudy (anonymous):

is this \[\sum\frac{x^{n+1}}{(n+1)!}\]?

OpenStudy (anonymous):

what am i saifoo now?

OpenStudy (anonymous):

this is \[e^x\]

OpenStudy (anonymous):

or maybe it is \[e^x-1\] because you are starting at n + 1 instead of n

OpenStudy (anonymous):

who is this?